Science Report Nov 15th

Science Report 15 November 2019

SOL: 5

Crew 215 – Expedition Boomerang Crew

Submitted by Science Officer Andrew Wheeler

Science Dome Operations: More micrometeorite samples were collected in the morning and

processed in the afternoon. Further gypsum samples were ground in preparation for the

cement making experiments to be carried out later in the rotation.

Green Hab: Green Hab operations have continued. See Green Hab report.

RAM Operations: N/A

EVA: The fifth EVA for science saw us returning to the MM5 (518462E 4248949N) site

where 24 micrometeorite samples were collected. 64% of the grid has now been sampled.

See EVA Report.

EVA Suit Maintenance: Nominal. See Operations Report.

Psycho-social Stressors Monitoring: Data collection continues through the hexoskin

monitoring vests. The simulated emergency retrieval of the engineer was carried out

during the engineering EVA and the performance of all participants is being evaluated.

In addition, indicated modifications to the airlock that would make this activity

easier are being evaluated.

Additional activities: N/A.

Research Report – November 13th

Science Report 13 November 2019
SOL: 3
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: N/A
Green Hab: Green Hab operations have continued. See Green Hab report.
RAM Operations: N/A
EVA: The third EVA for science saw us returning to the White Moon location at the junction of Brahe Hwy and Cow dung Road (517790E 4254750N) to collect more gypsum samples for ISRU cement experiments. It was noted that, at this location out of sim, at least one vehicle has been driving through the exact locations of our previous sampling efforts. Fortunately, this has no effect on our sampling, but it does suggest that our activities are been monitored extremely closely. See EVA Report.
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin monitoring vests. In addition, equipment for the retrieval of an injured marstronaut was inventoried prior to a rehearsal for an actual event. Subject to approval, a simulated emergency retrieval of the engineer will be held 14Nov2019 at the conclusion of his engineering EVA.
Additional activities: Nil.

Science Report – November 14th

Science Report 14 November 2019
SOL: 4
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: Gypsum samples retrieved were cleaned and weighed before
commencing grinding prior to experimentation. These samples were joined by
micrometeorite samples which were processed and packaged for shipment.
Green Hab: Green Hab operations have continued. See Green Hab report.
RAM Operations: N/A
EVA: The fourth EVA for science saw us returning to the MM5 (518462E 4248949N) site
where 30 micrometeorite samples were collected. 40% of the grid has now been sampled.
See EVA Report.
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin
monitoring vests. The simulated emergency retrieval of the engineer has been postponed,
subject to approval, to 15Nov2019 at the conclusion of his engineering EVA.
Additional activities: Hab operations were disrupted due to ongoing construction.

Research Report – November 12th

Science Report 12 November 2019
SOL: 2
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: The first 10 samples from the new micrometeorite sampling grid at MM5 (518462E 4248949N) were received, processed and prepared for dispatch. Previously collected gypsum samples were retrieved from storage in preparation for the ISRU cement/concrete manufacture experiments.
Green Hab: Green Hab operations have continued. See Green Hab report.
RAM Operations: N/A
EVA: The second EVA for science saw us returning to MM5 (518462E 4248949N) to complete setting out the micrometeorite sampling grid and collect the first row of 10 samples. See EVA Report.
EVA Suit Maintenance: Suit 10 has had the blown fuse replaced and is ready to be returned to use. Suit 2 is still being evaluated. See Operations Report.
Additional activities: Acoustic volumes were recorded. Habitat layout was evaluated. Psycho-social stressors continued to be monitored.

Science Report – Nov 11th

Science Report 11 November 2019
SOL: 1
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Dome Operations: Activity in the science dome included preparing for the deployment of the micrometeorite sampling grid and labeling the sample bags.

Green Hab: Green Hab operations have continued. See Green Hab report.

RAM Operations: The RAM Dome has been activated and has been accessed for suit batteries and fuses for Suit 10 and Suit 2. See the Operations report.

EVA: The first EVA for science involved a geological reconnaissance to the southwest of the hab where two locations on wash areas were scouted for the continuing sampling for micrometeorites. These sites – MM5 (518462E 4248949N) and MM6 (519990E 4248036N) appear to be seasonal alluvial fan deposits and the sampling grid has been established at MM5. See EVA Report.

EVA Suit Maintenance: The crew engineer has evaluated Suit 10 and Suit 2 and has identified the likely faults. See Operations Report.
Additional activities: Nil.

Science Report – Nov 8th

Crew 214 Science Report 8 November 2019

SOL: 12

Crew 214 – Expedition Boomerang Crew

Submitted by Science Officer Andrew Wheeler

Science Operations: No scheduled experimentation was carried out during the day. All micrometeorite samples collected during the rotation have been packaged and are awaiting dispatch to the Principle Investigator. All unused gypsum and regolith samples have been placed in storage awaiting use by Crew 215.

Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. Plants were watered by the Outpost crew. Additional seeds have been provided for sowing of further varieties.

Science Dome: Cleaned and tidied for Crew 215.

Closed Loop Food Waste Study: The final day of this study was concluded with the day’s waste for Sol 12 being: 28 grams; consisting of 99 calories, 2g fat, >1g saturated fat, 259mg sodium, 17g carbohydrates, 3g fibre, 3g sugar, 7g protein. For the entire 13-day rotation in sim, over 880g of food and nearly 4000 calories were wasted at a rate of approximately 7g of food per person per day. Carbohydrates contributed 63% of total calories wasted with fat and protein contributing 20% and 17% respectively.

EVA: No EVAs were performed.

EVA Suit Maintenance: Suit 10 and Suit 5 (helmet) remain unusable.

Additional Activities: N/A

Science Report Nov 8th

Crew 214 Science Report 8 November 2019

SOL: 12

Crew 214 – Expedition Boomerang Crew

Submitted by Science Officer Andrew Wheeler

Science Operations: No scheduled experimentation was carried out during the day. All micrometeorite samples collected during the rotation have been packaged and are awaiting dispatch to the Principle Investigator. All unused gypsum and regolith samples have been placed in storage awaiting use by Crew 215.

Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. Plants were watered by the Outpost crew. Additional seeds have been provided for sowing of further varieties.

Science Dome: Cleaned and tidied for Crew 215.

Closed Loop Food Waste Study: The final day of this study was concluded with the day’s waste for Sol 12 being: 28 grams; consisting of 99 calories, 2g fat, >1g saturated fat, 259mg sodium, 17g carbohydrates, 3g fibre, 3g sugar, 7g protein. For the entire 13-day rotation in sim, over 880g of food and nearly 4000 calories were was wasted at a rate of approximately 7g of food per person per day. Carbohydrates contributed 63% of total calories wasted with fat and protein contributing 20% and and 17% respectively.

EVA: No EVAs were performed.

EVA Suit Maintenance: Suit 10 and Suit 5 (helmet) remain unusable.

Additional Activities: N/A

Science Report – Nov 6th

Crew 214 Science Report 6 November 2019
SOL: 10
Crew 214 Ð Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: The seventh scientific EVA for the mission saw us return to Tank Wash #2 to collect 23 samples from the micrometeorite grid. Four of these had been compromised and it remains to be seen how this will affect the study. 64% of the plots have now been sampled.

With the disappointing failure of the regolith heater prior to arrival at MDRS, a modified plan to demonstrate water extraction was devised. Instead of an in situ test of the device, the gypsum and gypsiferous soil samples collected yesterday would undergo the process in the Science Lab. After documenting the samples, 6 small subsamples of ground gypsum were heated. Within minutes of each subsample being heated, water vapour was observed and video recorded. Further testing of the gypsiferous soil will be attempted tomorrow.

Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. The plants were watered by the Outpost crew.
Science Dome: Samples returned from yesterday’s EVA were processed and experimental procedures commenced.

Closed Loop Food Waste Study: The day’s waste for Sol 10 is: 51 grams; consisting of 175 calories, 4g fat, 2g saturated fat, 328mg sodium, 26g carbohydrates, 3g fibre, 8g sugar, 10g protein.

EVA: Preparations for finalizing the micrometeorite sampling experiment were undertaken.

EVA Suit Maintenance: Nominal

Additional Activities: Photographs were taken of the compromised micrometeorite sampling grid.

Science Report – November 05th

Crew 214 Science Report 5 November 2019
SOL: 9
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Operations: The sixth scientific EVA for the mission saw us travel to White Moon (517750E 4254520N) at the junction of Brahe Highway and Cow Dung Road in search of hydrated minerals. Immediately on arrival, gypsum crystals (CaSO4.2H2O) were observed protruding from the regolith and we were able to collect two soil samples and 9 rock samples. From these we intend to demonstrate the liberation of water through heating from both the crystals and the soil.
On our return to the the hab, we called into the Tank Wash #2 site and collected an additional 3 samples for the micrometeorite search. This is 43% of the planned samples. Unfortunately, someone has visited the grid, trampled through a number of plots and compromised the integrity of the experiment.
Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. The plants were watered by the Outpost crew.
Science Dome: Samples returned from the morning’s EVA are awaiting processing before undergoing experimental procedures within the science dome.
Closed Loop Food Waste Study: The day’s waste for Sol 9 is: 25 grams; consisting of 127 calories, 1g fat, <1g saturated fat, 69mg sodium, 70g carbohydrates, <1g fibre, 2g sugar, 5g protein.
EVA: Preparations for finalizing the micrometeorite sampling experiment were undertaken.
EVA Suit Maintenance: Nominal
Additional Activities: Photographs were taken of the compromised micrometeorite sampling grid.

Science Report – November 4th

Crew 214 Science Report 4 November 2019
SOL: 8
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: The fifth scientific EVA for the mission saw us returning to Tank Wash #2 and collecting 28 additional samples for the micrometeorite search. This is 40% of the planned samples.

Green Hab: Following the commencement of planting within the green hab, access through the tunnel between the main hab and the green hab has been opened. The plants were watered with approximately 3.5 litres of water, with a remaining 4.5 litres remaining in the watering can. The plant soil appeared fairly moist, so it is suggested crew refrain from over watering the new plants.

Science Dome: Samples returned from the morning’s EVA have now been processed within the science dome and limited observations of material recovered through the binocular microscope have been carried out.

Closed Loop Food Waste Study: The day’s waste for Sol 8 is: 51 grams; consisting of 159 calories, 3g fat, >1g saturated fat, 199mg sodium, 42g carbohydrates, 3g fibre, 13g sugar, 6g protein. Extra food supplies were received and added to the baseline stock database.

EVA: Preparations for an EVA to White Moon at the junction between Brahe Highway and Cow Dung Road have been carried out.

EVA Suit Maintenance: Images of helmet 5 have been forwarded to CapCom detailing the partial failure, the battery to Suit 10 has been examined and it is the suit itself that is faulty.

Additional Activities: Nil.